# Rubric Is All You Need: Enhancing LLM-based Code Evaluation With Question-Specific Rubrics

### v1

@@ 4.2. Complete Rubric Evaluation (CRE)
While the LLM performs only logical evaluation, syntax evaluation is conducted separately using a deterministic compiler-equipped agent. The student code is executed by the agent using a compiler via a system call in a Python script. The compiler returns a syntax assessment, and the agent uses a penalty-based system to assign syntax marks accordingly. For instance, with a maximum of 5 syntax marks and a penalty of 0.5, a code snippet with 5 syntax errors is awarded 2.5 marks. The final student marks are the sum of logical and syntactical scores. Prompts used for instructing LLMs are shared in Appendix A
@@ 4.3. Pointwise Rubric Evaluation (PRE)
Pointwise Rubric Evaluation (PRE) is a modified version of CRE. Instead of evaluating the entire rubric at once, the PRE LLM grader agent assesses the student code based on a single rubric point at a time. As shown in Figure 9, PRE Grader takes as input the problem statement, student code, and a single rubric point for evaluation. It returns a JSON dictionary containing the rubric point as the key and the assigned LLM marks as the value. These results are stored in a JSON file for later computations.
PRE is resource-intensive due to multiple API calls per student solution. An LLM call is executed in order to evaluate each point in the rubric, consuming tokens and time.
Figure 9. Pointwise Rubric Evaluation (PRE)
@@ 5.1. LLM selection
In order to maintain a balance between LLM accuracy and token costs, we used OpenAI’s GPT-4o-mini model to conduct our evaluations. The model has been shown to be remarkably accurate and consistent while maintaining significantly lower API costs versus OpenAI’s other models, including GPT-4o, o1 or o1-mini. We also used Anthropic AI’s Claude 3.7 Sonnet model with Extended Thinking deactivated.
@@ 5.2.1. Scaling
With varying scales and degrees of rubrics, the scale at which grades are awarded is different for every technique. Here, scale refers to the maximum marks or score that can be awarded for the evaluation technique (e.g. 35 for the OOP question-specific rubric, or 4 for the question-agnostic rubric). For the purposes of understanding, we shall refer to the Ground Truth values as the Base Dataset (B) and the LLM evaluations as the Experiment Dataset (E). $B[i]$ and $E[i]$ denote the base and experimental marks obtained by the $i^{th}$ student, where $0\leq i<N$. To assess the correspondence of an Experiment Dataset E that lies on a different scale versus B, we scale E linearly as
$E_{s}[i]={E[i]}\times\frac{R}{R_{E}},\forall 0\leq i<N$
where $R_{E}$ (e.g. 4 for the question-agnostic rubric) is the scale of the experimental rubric, $R$ is the scale of the question-specific rubric (also the scale of ground-truth values, e.g. 35 for the OOP dataset), $N$ is the number of data points in the dataset and $E_{s}$ is the scaled experimental dataset. For techniques evaluated method-by-method on the OOP dataset using question-agnostic rubrics, we scaled $E$ on the method level as
$E_{s}[i]=\sum_{m=1}^{M}{(E[i][m]\times\frac{R[m]}{R_{E}})},\forall 0\leq i<N$
where $m$ is the method being evaluated, $M$ is the number of methods present in the problem statement and $R[m]$ are the maximum number of marks that can be awarded for the method in the question-specific rubric. The total marks for a student would therefore be the sum of marks awarded for each method, scaled up to match the base rubric.
@@ 5.2.2. Binning
Before we obtain Cohen-Kappa scores on continuous datasets, we first classify the data into 5 bins. The bins are decided based on the ranks of the data points in the sorted datasets. For instance, the dataset $[3,2,4,3,1,9,6,7]$ when classified into 3 bins is graded as $[0,0,1,1,0,2,1,2]$, where 0, 1 and 2 are the three possible grades, assuming uniform grading.
@@ 5.3. Evaluation Metrics
In order to assess the performance of question-specific rubrics versus question-agnostic rubrics on the proposed datasets, we shall use the following metrics:
  • Pearson Correlation Coefficient (PCC, $r$) (sci, 2025): PCC measures linearity between two datasets. A value of magnitude 1 indicates perfect linearity between the sets. Positive values indicate an increasing nature in $E$ as $B$ increases, while negative values indicate a decreasing nature.
  • Spearman Rank Correlation Coefficient ($r_{s}$) (sci, 2025): Spearman Correlation measures the degree of monotonicity between the two sets. Similar to the Pearson coefficient, an absolute value of 1 indicates perfect monotonicity, while positive and negative values denote increasing and decreasing trends in $E$ while $B$ increases.
  • Kendall-Tau Correlation Coefficient ($\tau_{b}$) (sci, 2025): Also called the Kendall Rank Correlation Coefficient, it measures ordinal association between two quantities. It is a measure of rank correlation, i.e., it denotes the similarity of the orderings of data points when ranked. Values of positive or negative 1 indicate perfect association, while the sign of the coefficient denotes an increasing or decreasing trend.
  • Leniency (Mean Normalized Error, $l_{n}$): The leniency between two datasets is defined as
    $l_{n}=\frac{\sum_{i=0}^{N-1}\hat{E}[i]-\hat{B}[i]}{N}$
    where $\hat{E}[i]$ and $\hat{B}[i]$ are the experimental and base scores normalized to 1. Leniency is a measure of how strict the evaluation system is for a rubric. Leniency is relative, and requires base data against which the metric is computed. A perfectly lenient grader would assign R marks where every base data point is 0, achieving a leniency score of 1. A perfectly strict grader would assign 0 marks where every base data point is R, achieving a leniency score of -1. If we consider a target average deviation of 10%, the leniency to aim for would be within $\pm$0.1.
  • Intraclass Correlation Coefficient (ICC) (pin, 2024): The ICC is used to assess the consistency of measurements made by multiple observers that measure the same quantity. We measure the ICC of the llm-based evaluation techniques graders against the base dataset $B$ in order to estimate how well the experimental scores match the truth. We use the following 3 ICC scores
    • One-way random effects (ICC1): each subject is measured by a different set of k randomly selected raters;
    • Two-way random (ICC2): k raters are randomly selected, then, each subject is measured by the same set of k raters;
    • Two-way mixed (ICC3): k fixed raters are defined. Each subject is measured by the k raters.
  • Cohen-Kappa w/ Binning (skl, 2024) ($\kappa_{B}$): As described in the previous section, $\kappa_{B}$ is the Cohen-Kappa score obtained after binning the data points into a predetermined number of bins (5 in our case). This is done because Cohen-Kappa measures correspondence between graded scores. Binning allows us to emulate relative grading to an extent.
@@ 6.1. Techniques Used
  • Five Point Marking (FPM) (Phung et al., 2023): Five Point Marking is a question-agnostic LLM-based rubric marking technique. The student code is evaluated based on the following points:
@@
For a total score of 100. The scores are scaled down to 35 for comparison with ground truth values.
@@ 6.2. Findings
Tables 2 and 3 provide an overview of the performance of all implemented techniques on the OOP and DSA datasets. Table 2 provides an overview of the evaluation performed on the OOP dataset, with results from EME, CRE and PRE, while Table 3 provides an overview of the results obtained for the DSA dataset. We evaluate the question-agnostic versus question-specific rubrics on EME.
@@
Question-Agnostic Rubric
@@
Question-Specific Rubric
@@
Question-Agnostic Rubric
@@
Question-Specific Rubric
@@
Table 3. Results for Question-Agnostic and Question-Specific rubrics on the DSA dataset
Within the scope of this paper, we have discussed building rubrics, and the differences between question-specific and question-agnostic rubrics. We have also introduced evaluation techniques for question-specific rubrics, and compared the performance of LLM grading with question-specific rubrics versus question-agnostic rubrics.
While PRE is instructed to evaluate only logical correctness, it is observed to be more strict and detail-oriented than CRE. PRE consistently awards lower marks compared to CRE, leading to lower correlation metrics, as human grading of the OOP dataset was relatively lenient. The primary reason for this is that PRE, when given a single rubric point, applies stricter criteria. Since the LLM evaluates only one aspect at a time, it tends to assign zero marks unless the code fully meets the given rubric criteria. PRE is suitable for use cases requiring stringent evaluation and strict rubric adherence.
For EME, we observe the following:
@@
  • Ensemble Size: The effectiveness of the method varies with ensemble size. Initially, a significant increase in correlation coefficients is observed as ensemble size decreases. This trend continues until the ensemble reaches an optimal range—typically around three to four models when employing high-parameter LLMs, such as GPT-4 and Claude 3. Beyond this threshold, performance gains tend to plateau, suggesting diminishing returns with further increase in ensemble size.
We observe clearly that question-specific rubric based evaluation techniques outperform all other techniques with respect to the metrics we have used. EME (QS) outperforms EME (QA) on DSA datasets, achieving high correlation, and performing the best out of all techniques in terms of correlation coefficients, leniency, ICC and $\kappa_{B}$. 3
CRE scores consistently high for all correlation scores, and takes the lead with Pearson Correlation as well as with ICC3 and $\kappa_{B}$. 2 PRE has a leniency of -0.329 against ground truth values. This means that on average, PRE scores $35\times 0.329=11.515$ marks lower than the ground truth. In comparison, CRE has a leniency of 0.081, which in magnitude is considerably lower than that of PRE. This corresponds with our hypothesis that LLMs given a single rubric point will be considerably stricter with the evaluation of the student code.
6.3. Variations in LLM Grading Performance
6.3.1. Dataset
: Our techniques consistently yield better performance on the OOP dataset, achieving higher scores across various evaluation metrics. We observe that QA and QS techniques perform relatively well with the OOP dataset. This can be attributed to the nature of the dataset, which consists of questions that are relatively straightforward to implement, with minimal variation in algorithmic approaches or implementation strategies across different solutions. These questions primarily pose an implementation challenge rather than requiring deep logical reasoning. As a result, advanced language models like GPT-4 and Claude 3.7 perform well even without a predefined rubric. In contrast, their performance declines on the DSA dataset, which demands a deeper understanding of algorithms and their underlying logic. The results indicate that while LLMs effectively evaluate solutions in the OOP dataset, they struggle with the greater algorithmic diversity and complexity present in the DSA dataset, especially when a question-specific rubric is absent.
6.3.2. Presence of LLM Grader
: We observe that LLM-based techniques outperform CodeBERTScore significantly and impressively. CodeBERTScore obtains weak scores across all evaluation metrics. LLMs are vastly more versatile when it comes to understanding context, as well as the varying nature of student approaches to one particular problem.
6.3.3. Presence of Rubric
: We observe that both question-agnostic and question-specific rubric techniques outperform no-rubric techniques. While CodeJudge achieves respectable correlation and $\kappa_{B}$ scores, providing a rubric gives the LLM grader an anchor around which to evaluate or grade the code, thereby increasing performance significantly when a rubric is provided.
6.3.4. Specificity of Rubric
: We observe that question-specific rubric techniques outperform question-agnostic rubric techniques in the DSA dataset, while both question-specific and question-agnostic techniques score high within the OOP dataset. We provided the LLM grader the exact rubric that human graders were given, and observe that the LLM grader is able to follow the marking scheme with a satisfying degree of correctness, as evidenced by the consistently high correlation scores, and the observed leniency scores within the range $0.0_{\pm{0.1}}$.
NOTE: We exclude PRE from this comparison for the reasons described next.
6.3.5. Manner of Feeding
: We observe that the manner in which a rubric is presented to a large language model (LLM) can significantly influence the grading outcomes. In the EME framework, we provide a model solution as part of the prompt. This inclusion likely accounts for the comparable performance of EME(QA) and EME(QS), as object-oriented programming (OOP) student implementations tend to closely align with the provided model solution. Conversely, in the case of data structures and algorithms (DSA) evaluation, the model solution has limited utility, as student implementations exhibit substantial variation. Within EME (QS) and CRE, we feed the LLM the entire rubric at once, after which the LLM grader performs grading in a single LLM call. Within PRE, each rubric evaluation point is fed one-by-one. When fed evaluation points one by one, the LLM is remarkably strict in grading, as evidenced by the $l_{n}$ score of -0.329. We therefore make the claim that to emulate grading on a human-like scale, the LLM grader should be fed the entire rubric as an evaluation context.

### v2

@@ 4.2. Complete Rubric Evaluation (CRE)
While the LLM performs only logical evaluation, syntax evaluation is conducted separately using a deterministic compiler-equipped agent. The student code is executed by the agent using a compiler via a system call in a Python script. The compiler returns a syntax assessment, and the agent uses a penalty-based system to assign syntax marks accordingly. For instance, with a maximum of 5 syntax marks and a penalty of 0.5, a code snippet with 5 syntax errors is awarded 2.5 marks. The final student marks are the sum of logical and syntactical scores. Prompts used for instructing LLMs are shared in Appendix A.
@@ 4.3. Pointwise Rubric Evaluation (PRE)
Pointwise Rubric Evaluation (PRE) is a modified version of CRE. Instead of evaluating the entire rubric at once, the PRE LLM grader agent assesses the student code based on a single rubric point at a time. PRE Grader takes as input the problem statement, student code, and a single rubric point for evaluation. It returns a JSON dictionary containing the rubric point as the key and the assigned LLM marks as the value. These results are stored in a JSON file for later computations. PRE is resource-intensive due to multiple API calls per student solution. An LLM call is executed in order to evaluate each point in the rubric, consuming tokens and time.
@@ 5.1. LLM selection
In order to maintain a balance between LLM accuracy and token costs, we used OpenAI’s GPT-4o-mini model to conduct our evaluations. We also used Anthropic AI’s Claude 3.7 Sonnet model with Extended Thinking deactivated.
@@ 5.2.1. Scaling
With varying scales and degrees of rubrics, the scale at which grades are awarded is different for every technique. Here, scale refers to the maximum marks or score that can be awarded for the evaluation technique . In the OOP dataset, the question-specific rubric awards up to 35 points, whereas the question-agnostic rubric compresses the same outcome space into a holistic 5-point scale. Moreover, when evaluated method-by-method, a question-agnostic rubric evaluates each method on the same 0-4 scale, whereas in the original rubric, each method holds different weights in terms of maximum marks that can be awarded for that method. Therefore, in order to make quantitative comparisons between the two datasets, we scale the grades awarded by evaluation on the question-agnostic rubric to match the 35 mark question-specific rubric design. For the purposes of understanding, we shall refer to the Ground Truth values as the Base Dataset (B) and the LLM evaluations as the Experiment Dataset (E). $B$ and $E$ contain $N$ data points, one for each student being included in the assessment. $B[i]$ and $E[i]$ denote the base and experimental marks obtained by the $i^{th}$ student, where $0\leq i<N$. We use the terms ”Base” and ”Experiment” dataset with the belief that the metrics and techniques used for comparing the grading performances of an LLM grader vs a Human grader can be used to confirm the degree of affirmation between any two graders regardless of their nature, where the Experiment Grader is compared to a Base Grader in order to assess agreement. We note that for evaluation on a method level (as seen in PRE), the experiment dataset $E$ is two dimensional, where for each student $i$, $E[i]$ contains $M$ data points, where $M$ is the number of methods being evaluated, and $E[i][m]$ denotes the marks obtained by student $i$ for the method $m$. To compute metrics for an Experiment Dataset E that lies on a different scale versus B, we scale E linearly as
$E_{s}[i]=E[i]\times\frac{R}{R_{E}},\qquad 0\leq i<N$
where $R_{E}$ (e.g. 4 for the question-agnostic rubric) is the scale of the rubric being used for the evaluation of experiment marks, $R$ is the scale of the question-specific rubric (also the scale of ground-truth/base values, e.g. 35 for the OOP dataset), $N$ is the number of data points in the dataset, and $E_{s}$ is the scaled experimental dataset.
For techniques evaluated method-by-method on the OOP dataset using question-agnostic rubrics, we scaled $E$ on the method level as
$E_{s}[i]=\sum_{m=1}^{M}\bigl{(}E[i][m]\times\tfrac{R[m]}{R_{E}}\bigr{)},\qquad 0\leq i<N$
where $m$ is the method being evaluated, $M$ is the number of methods present in the problem statement, and $R[m]$ is the maximum number of marks that can be awarded for method $m$ in the question-specific rubric. The total marks for a student are therefore the sum of marks awarded for each method, scaled up to match the base rubric.
The scaled experiment dataset $E_{s}$ therefore eliminates the dimensionality issue presented by evaluating the experiment dataset $E$ on a method level, by reducing it to a one dimensional dataset.
@@ 5.2.2. Binning
Before we obtain Cohen–Kappa scores on continuous datasets, we first classify the data into 5 bins. The bins are decided based on the ranks of the data points in the sorted datasets. For instance, the dataset $[3,2,4,3,1,9,6,7]$ when classified into 3 bins is graded as $[0,0,1,1,0,2,1,2]$, where 0, 1 and 2 are the three possible grades, assuming uniform grading. Concretely, each scaled score list is sorted and partitioned at the 20th, 40th, 60th and 80th percentiles, producing five ordinal intervals of roughly equal size. Marks falling below a cut-off receive the lower bin label (0–4). We adopt five bins because (i) it parallels the familiar A–E grading scale used in educational studies, (ii) it maintains adequate cell counts for stable $\kappa$ estimates with $N\!\approx\!100$, and (iii) sensitivity checks with 4 and 6 bins shifted $\kappa$ by no more than 0.02. We acknowledge that any discretisation can distort near-threshold scores, yet the empirical drift observed here is minor and does not affect qualitative conclusions.
@@ 5.3. Evaluation Metrics
We assess each rubric–LLM pair with seven complementary statistics that together expose ordering, bias and exact‐score agreement. Although several of the following statistics are mathematically related, no single number captures all facets of grading quality. These seven metrics can be categorized into three broad dimensions:
(1) Relative ordering of students (correlations) Rank correlations:
Pearson Correlation Coefficient (PCC, $r$) (sci, 2025) captures linear alignment and is most sensitive to large outliers. Spearman Rank Correlation Coefficient ($r_{s}$) (sci, 2025) tests whether the overall ranking is preserved regardless of spacing, while Kendall-Tau Correlation Coefficient ($\tau_{b}$) (sci, 2025) provides a tie‐aware alternative that remains stable on small samples. High cross-correlation among some measures (e.g. Pearson vs. Spearman) signals convergent validity, while the divergences highlight specific failure modes such as monotonic but non-linear trends or category-level disagreement after binning—that would be invisible in a single score.
(2) Agreement coefficients:
Intraclass Correlation Coefficient (ICC) (pin, 2024) comes in three flavours: ICC1 (one‑way random), ICC2 (two‑way random) and ICC3 (two‑way mixed). These differ in how they treat the raters (randomly drawn versus fixed). All express absolute agreement in the original mark scale. Cohen-Kappa ($\kappa_{B}$) (skl, 2024) treats scores as categories (bins), correcting for chance agreement; useful when instructors ultimately issue letter grades.
(3) Absolute bias or strictness:
Leniency (Mean Normalized Error, $l_{n}$)) between two datasets is defined as
$l_{n}=\frac{\sum_{i=0}^{N-1}(\hat{E}[i]-\hat{B}[i])}{N},$
where $\hat{E}[i]$ and $\hat{B}[i]$ are the experimental and base scores normalised to 1. Leniency is a measure of how strict the evaluation system is for a rubric. Leniency is relative, and requires base data against which the metric is computed. A perfectly lenient grader would assign $R$ marks where every base data point is 0, achieving a leniency score of 1. A perfectly strict grader would assign 0 marks where every base data point is $R$, achieving a leniency score of –1. If we consider a target average deviation of 10%, the leniency to aim for would be within ±$\pm 0.1$.
Why Leniency matters?
Correlation metrics (Pearson, Spearman, Kendall) reward graders that preserve the student ranking even when every mark is shifted by a constant amount. Leniency exposes that uniform bias: a positive $l_{n}$ signals generosity, a negative $l_{n}$ strictness. Considering both together separates disagreements caused by level shifts (bias) from those caused by rank dispersion. In practice we sometimes observe $r>0.9$ while $|l_{n}|$$\pm 0.15$, indicating that a simple post‑hoc rescaling (not rubric redesign) can reconcile the two graders.
@@ 6.1. Techniques Used
  • Five Point Marking (FPM) (Phung et al., 2023): Five Point Marking is a question-agnostic LLM-based rubric marking technique. The student code is evaluated out of 100 split into the following categories:
@@
The scores are scaled down to 35 for comparison with ground truth values.
@@ 6.2. Findings
Tables 2 and 3 provide an overview of the performance of all implemented techniques on the OOP and DSA datasets.
@@
Question-Agnostic (QA) Rubric
@@
Question-Specific (QS) Rubric
@@
Question-Agnostic (QA) Rubric
@@
Question-Specific (QS) Rubric
@@
Table 3. Results obtained by various techniques on the DSA dataset
6.2.1. Overall Observations (Table 2 and Table 3)
  • Presence of LLM Grader: We observe that LLM-based techniques outperform CodeBERTScore significantly and impressively. CodeBERTScore obtains weak scores across all evaluation metrics. LLMs are vastly more versatile when it comes to understanding context, as well as the varying nature of student approaches to one particular problem.
  • Presence of Rubric: We observe that both question-agnostic and question-specific rubric techniques outperform no-rubric technique (CodeJudge). While CodeJudge achieves respectable correlation and $\kappa_{B}$ scores, providing a rubric gives the LLM grader an anchor around which to evaluate or grade the code, thereby increasing performance significantly when a rubric is provided
6.2.2. Specific Observations from Results on DSA Dataset (Table 3)
[…]
  • Type of Question-Agnostic rubric: Using EME with a QA rubric provides significantly better correlation scores versus the FPM technique. The FPM technique attempts to grade the solution over 5 predefined marking points. An LLM when provided with only a problem statement and a solution code snippet is unable to grade accurately based on these 5 marking points, thereby yielding lower scores. In contrast, it performs better when asked to gauge the degree of correctness of the entire code snippet directly.
6.2.3. Specific Observations from OOP Results (Table 2)
  • Both question-specific and question-agnostic rubrics achieve comparable results for EME: We see comparable performance for question-specific (EMA(QS)) and question-agnostic (EMA(QA)) rubric-based approaches across all metrics. As the OOP dataset contains homogeneous implementation-oriented questions, the results indicate that such questions are graded reliably even with question-agnostic rubrics.
  • Strictness depends on prompt granularity: As shown in Table 2, PRE (feeding one rubric point at a time) slashes average scores by $11.5/35$ marks (leniency $=-0.329$), whereas CRE (feeding the whole rubric) hovers near human leniency $(0.081)$.
  • Why PRE is harsher than CRE? Single-criterion prompts force the model to assign zero unless the exact logic is present, whereas whole-rubric prompts let it award partial credit, mirroring human evaluator behaviour. PRE seems to be more suitable for use cases requiring stringent evaluation and strict rubric adherence.
6.2.4. EME Performance
We conducted an additional set of experiments to understand the impact of model parameter size and ensemble size for EME.
@@ 6.2.4. EME Performance
  • Ensemble Size: The effectiveness of the method varies with ensemble size. Initially, a significant increase in correlation coefficients is observed as ensemble size increases. This trend continues until the ensemble reaches an optimal range—typically around three to four models when employing high-parameter LLMs, such as GPT-4 and Claude 3. Beyond this threshold, performance gains tend to plateau, suggesting diminishing returns with further increase in ensemble size.
7. Discussion
7.1. Implications for Instructors and Teaching Assistants
For instructors and teaching assistants, the three graders (CRE, PRE and EME) can be combined in a staged workflow. A lightweight CRE pass may quickly triage submissions, flagging clearly strong or weak attempts, whereas stricter PRE scoring (or a brief human check) can be reserved for borderline cases. In settings where false positives carry a high cost, such as summative examinations, PRE (strict) may offer additional reassurance, trading a slight reduction in overall agreement for a zero‑tolerance stance on partial logic. Routine coursework and large MOOC cohorts, by contrast, often prioritise speed and cost. For such scenarios, CRE seems to be a reasonable cost-effective choice. Classes that attract highly diverse algorithmic solutions (for example, DSA) could find value in EME (QS), which attained the highest ICC3 in our DSA benchmark, albeit at the price of a three‑model ensemble. We also note that for larger, easier and more direct problems, as seen in the OOP dataset, a simple question-agnostic rubric with method-wise evaluation produces near-human-like grading and is on par with evaluation using question-specific rubrics. It does, however, rely on multiple LLM calls (once for each method) rather than a single LLM call with a large rubric. Developing a question-specific rubric may therefore save token costs during evaluation.
The JSON traces emitted by all three techniques may also feed into analytics dashboards. Aggregating the most frequently missed rubric criteria can highlight common misconceptions, enabling teaching teams to design targeted recitations, micro‑lectures, or discussion‑board posts that address exactly those weak spots.
A minimal integration path could involve: (i) translating an existing rubric into the key–value format accepted by the graders, (ii) calibrating Leniency or ensemble size on roughly ten pilot submissions until the mean‑normalised error falls within ±$\pm 0.1$, (iii) inserting the grader call into the continuous‑integration script that already compiles and tests student code, and (iv) offering reviewers an interface that surfaces rubric items with confidence below 0.8 for optional human override.
7.2. Guidance for Students
Rubric‑aligned feedback provides criterion‑level transparency: each comment is anchored to a specific step (e.g., Step 6 – create a Player object), helping learners understand why marks were lost instead of receiving a generic wrong output notice. Encouraging students to summarise the feedback in a simple worksheet (criterion missed, probable cause, planned fix) may promote systematic debugging. Learners might further consolidate gains by submitting a short reflection after resubmission detailing which rubric elements they have mastered, which remain challenging, and what strategies (additional unit tests, peer review, etc.) they will employ next time. Such reflective practice aligns with accepted principles of self‑regulated learning.